OpenStudy (really):

what is the lcd of 5 9 and 7 12

OpenStudy (anonymous):

\[5\times 9\times 7\times 4\]

OpenStudy (anonymous):

i hope the question actually asked "what is the least common MULTIPLE" not lcd you have no denominators there, just whole numbers

OpenStudy (anonymous):

It would be 3. I hope this helps.

OpenStudy (anonymous):

I think he meant 5/9 and 7/12. :)

OpenStudy (anonymous):

9=1,3 12=1,2,3 The three is the least common denominator.

OpenStudy (anonymous):

Thanks for the medal!